OpenStudy (princeevee):

the variable t represents the temerature at 4 p.m which equation would best model the word problem below the temperature at the summit of mt mansfield dropped 17*F between 4 p.m and 11 p.m. the temp at 11 p.m was -11*f t-17=-11 11-17=t 17-t=-11 17+11=t

OpenStudy (defiance):

Hiii!!! again!

OpenStudy (defiance):

t = temp at 4 PM

OpenStudy (defiance):

If it as dropped 17F from 4 - 11pm

OpenStudy (defiance):

If we're at -11 F at 11PM

OpenStudy (defiance):

Then add 17 to -11.

OpenStudy (princeevee):

hello earth human again )=)=: D

OpenStudy (defiance):

Do you still need help?

OpenStudy (princeevee):

which is 6

OpenStudy (princeevee):

i guess

OpenStudy (defiance):

Yes, t is 6.

OpenStudy (princeevee):

but which is the right equation?

OpenStudy (defiance):

Okay.

OpenStudy (defiance):

The numbers are listed. With the acception of "t" Which we have found is 6. D. Makes no sense 17 + 11 = 28. So, incorrect 11 - 17 = -6. Therefore, incorrect.

OpenStudy (defiance):

t - 17 = -11 is correct, BUT. Pretend we don't now what T is. So that wouldn't help us solve it.

OpenStudy (defiance):

Oops

OpenStudy (defiance):

Sorry.

OpenStudy (defiance):

A is correct. I was thinking about C but i didnt read it right.

OpenStudy (defiance):

C. 17 - 6 = -11. No. Cause thats not true. Leaves you with A

OpenStudy (princeevee):

so a is the correct answer?

OpenStudy (defiance):

Your answer is A, cause it's the only one that is correct.

OpenStudy (defiance):

Yes!

OpenStudy (princeevee):

ok thanks earth human

OpenStudy (defiance):

Sorry about being confused for a moment.

OpenStudy (princeevee):

its ok

OpenStudy (defiance):

Very welcome!